5. A spherical snowball with an outer layer of ice melts, so that the radius of the snowball


decreases at the rate of 1/5 cm/sec. Find the rate at which the volume decreases when the


diameter is 50 cm.

1
Expert's answer
2021-12-21T01:49:02-0500

The formula for volume of a sphere is "V=\frac{4}{3} r^{3} \pi."

Differentiating with respect to t, time.

"\frac{d V}{d t}=4 r^{2}\left(\frac{d r}{d t}\right)"

The rate of change of the snowball is given by "\frac{d V}{d t}."

We know "\frac{d r}{d t}=-\frac{1}{5}."

We want to find the rate of change when "d=50\Rightarrow r=\frac{50}{2}=25."

Hence,

"\begin{aligned}

&\frac{d V}{d t}=4(25)^{2}(-\frac{1}{5}) \\

&\frac{d V}{d t}=-500

\end{aligned}"

Thus, the volume of the snowball is decreasing at a rate of "500 \frac{\mathrm{cm}^{3}}{\mathrm{sec}}."